Raith Status report 01052005:1230 hours System access delayed through Friday.

James Conway jwc at snf.stanford.edu
Wed Jan 5 13:51:21 PST 2005


Greetings Raith Users:

There will be a delay in the previously announced schedule 
<http://snf.stanford.edu/cgi-bin/ezmlm-cgi?2160:mss:1031:200412:khjpoomaipbapbcfmmld> 
to return the system to users access through Friday Noon.
Users with  reservations for RAITH on the CORAL schedule before Noon 
Friday should remove their reservations and reschedule for a later time.

Over the holiday we completed a 200 hour bake out of the System after 
the PM routine was performed just before the shut down. We replaced the 
Field Emission Gun tip, aperture plate, and performing many cleanings 
and inspection routines of various sub-systems on the RAITH.

Completed thus far were the following:

   1. Replaced the FE- Gun and multi-hole aperture plate with new parts.
   2. System Administration to both the LEO and RAITH computers,
      updating all windows patches and antivirus software definitions. 
      Swept the system of viruses and spy ware.  NO Viruses found.
      Backed up the system to HD.
   3. Load Lock and transfer arm and track maintenance as scheduled.
      Also tensioned the bands to this section
   4. Inspection and Cleaning of the System Chamber and Piezo control
      kinetic mounts. Check exchange positions for transfer arm.
   5. Pole Piece removed inspected and cleaned.
   6. Realigned the laser interferometer optimizing the laser signal to
      12 Volt Peak to Peak, an optimal value.
   7. Stage Motor Controller: Calibrated the X/Y Sin-Cosine control
      board and checked the motors for current draw values. Tested the
      motor controller power supply output voltage.
   8. Exchanged the loaner Water Chiller with our repaired unit.
   9. Performed an overnight bake out and a brief 'dry' run up of the
      FE- gun up to 20 keV without the EHT voltage applied to evaluate
      the cleanliness of the anode and aperture plates and FE-gun to
      evaluate arcing concerns.

The extended bake out of the system completed  1800 hours December 25, 
2004 and upon cooling the gun section reached 9.8 E -011 mBar.  Upon 
arrival yesterday I observed 1.8 E - 010 mbar vacuum.  This is wonderful 
vacuum level to start conditioning the FE-Gun and performing the service 
run up greatly reducing the chance of FE-Gun arcing during the service 
run up.

Yesterday afternoon we started the service and conditioning run up of 
the FE-Gun.  While one minor EHT trip event was encountered as we 
initially brought the Fe-Gun on line, the rest of the ramp to 1000 Volts 
went smoothly and we were able to maintain good vacuum.  We will be 
continuing this work, resuming this afternoon, and we plan to perform a 
very slow and cautious run up to 10 KeV level before we start alignments 
to the mechanical position of the FE-Gun, check out the E/O section of 
the system, and finally performing all the usual calibrations and 
adjustments of the gun, aperture(s), and stigmation centering functions. 
After ~24 hours of beam time at this level I hope to be able to qualify 
the system and hand off access to users late this week.

NOTE: There remain a number of issues still to be addressed to 
effectively eliminate all the issues we encountered in the last quarter 
of the year mainly related to image drift on image acquisition during 
Write Field alignment; leakage and alignment of the Raith beam Blanker; 
and elimination of the extra shot spots on reduced scan fields. The 
remain a few other nagging concerns related to Single Pixel Line 
resolution and circularity of Circle and dot shot features written with 
the system we hope will be eliminated by the FE-Gun exchange and PM of 
the SEM column.

I expect this work to take much of the week to complete but will offer 
some access to RAITH Champions scheduled on the system after Friday Noon 
and through the weekend.  But with the caveat that only those whom are 
willing to evaluate their results and promptly return to me their 
measurements will get access to the system until I have finished all 
outstanding action items.

FYI:  I am writing up a document that describes the PM routine task and 
will add this to my System Qualification Process Control Monitoring 
(PCM) testing documentation previously posted and also in RAITH Notebook 
No. 2 in the Lab.

Thank you for your support!

James Conway



Previous announcement WED DEC. 8, 2004: 
<http://snf.stanford.edu/cgi-bin/ezmlm-cgi?2160:mss:1031:200412:khjpoomaipbapbcfmmld>

Greetings RAITH Users:

We will be taking the system off line from Monday December 13, 2004  at 
10 00 AM  through January 5, 12 00 hours.
The coral schedule has been blocked out for this period. Please do not 
attempt to make reservation during this period.

We will make every effort to have the system back up completing all 
adjustments to the system, and have the system performance qualified, by 
NOON January 5, 2005 if all goes well through the PM, shutdown, start up 
and qualification testing period.

Should unforeseen circumstances cause additional delays I will be 
posting to this email discussion list the afternoon of January 4, 2005.  
Please stay tuned.

Thank you for your support!

James Conway
SNF


-------------- next part --------------
An HTML attachment was scrubbed...
URL: <http://snf.stanford.edu/pipermail/raith/attachments/20050105/071ad471/attachment.html>


More information about the raith mailing list